# CAF-I: A Collaborative Multi-Agent Framework for Enhanced Irony Detection with Large Language Models

### v1

@@ Abstract
Multi-Agent System, Irony Detection, Large Language Model
@@ 1 Introduction
The rapid development of social media has fostered increasingly diverse and semantically complex forms of textual expression (Kader et al., 2022) highlights the challenge of irony detection. Irony, a rhetorical device contrasting stated and intended meaning (Booth, 1974), is crucial for NLP systems as its presence alters textual interpretation, impacting applications like sentiment analysis and content moderation (Reyes et al., 2013). However, detection is complicated by context dependency, obscure authorial intent, and varied rhetorical strategies (Wallace et al., 2015).
Initial efforts used traditional machine learning with handcrafted features (Reyes et al., 2013), often failing to capture irony’s subtleties. Deep learning methods improved semantic representation but struggled with the implicit and contextual nature of ironic utterances (Ghosh & Veale, 2016; Baziotis et al., 2018). Consequently, the emergence of LLM has presented novel avenues, exhibiting substantial potential in irony detection, typically via fine-tuning or advanced prompt engineering techniques (Yao et al., 2024).
@@
Motivated by Chain-of-Thought (CoT) prompting (Wei et al., 2022) and multi-agent cooperation principles (Liu et al., 2023), this paper introduces CAF-I, a novel LLM-driven multi-agent framework to address existing limitations in irony detection by simulating human-like multidimensional analytical processes. CAF-I initially performs multi-perspective parallel analysis from three core angles: contextual consistency, deep semantic logic, and rhetorical function. Subsequently, an interactive collaborative optimization mechanism facilitates information sharing and debate among these perspectives. Finally, an integrated decision mechanism allows a central decision-making agent to consolidate these viewpoints and apply conditional optimization triggered by evaluative feedback, forming the paper’s primary contributions:
@@ 2.1 Irony Detection
Irony detection in NLP remains challenging due to its inherent complexity, context dependence, and contradictory nature. Early machine learning relied on hand-crafted features, including lexical cues (Davidov et al., 2010) and sentiment lexicons (Reyes et al., 2013), but inadequately captured irony’s subtleties. Deep learning techniques offered significant improvements: Word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) enabled richer semantic representations. Subsequently, Convolutional Neural Networks (Poria et al., 2016), recurrent neural networks, and particularly LSTM architectures (Zhang et al., 2016) learned hierarchical and sequential features, while Graph Neural Networks (Liang et al., 2022) captured structural dependencies to improve performance. Nonetheless, these deep learning approaches still faced difficulties with irony’s implicit meanings and complex rhetoric, motivating the exploration of more advanced models.
Recent advances in large language models have introduced prompt-based and zero-shot learning methods, allowing models like GPT-4o to interpret irony with minimal reliance on labeled data (Yao et al., 2024). By leveraging task-specific prompts, these models have demonstrated enhanced context-aware irony classification. However, existing LLM-based approaches still operate as single-model predictors, lacking the capacity for multi-perspective reasoning.
@@ 2.2 CoT Prompt Engineering
To address the limitations of direct predictions in complex reasoning tasks, CoT prompting emerged as a structured technique that enhances language models by guiding them through intermediate reasoning steps (Wei et al., 2022). Despite its effectiveness, CoT initially depended on manually created prompts, which constrained its adaptability. Subsequent innovations, such as Auto-CoT, addressed this by automating the creation of reasoning chains (Zhang et al., 2022). Further extensions, such as Tree-of-Thought (ToT), expanded reasoning by enabling the exploration of multiple pathways (Yao et al., 2024), while Graph-of-Thought (GoT) introduced a structured approach where reasoning steps are represented as interconnected nodes in a graph (Besta et al., 2024). These advancements directly influenced our multi-agent framework, where specialized agents leverage structured reasoning to collaboratively enhance irony detection.
@@ 2.3 Multi-agent Cooperation
Multi-agent frameworks effectively leverage collaborative interactions among specialized agents for complex tasks, with studies exploring forms like deliberation, structured debates, and dialogues. Debating frameworks, for instance, improve factual accuracy and solution diversity in complex reasoning (Du et al., 2023). Others, such as CAMEL (Li et al., 2023), use role-play to simulate nuanced human behaviors; MathChat (Wu et al., 2023b) employs structured dialogues for intricate tasks; and AutoGen (Wu et al., 2023a) shows multi-agent versatility through customizable structures. Inspired by these methodologies, our framework uses specialized LLM-based agents for contextual, semantic, and rhetorical analysis via CoT reasoning. A final decision agent integrates their outputs, enhancing irony detection’s robustness, accuracy, and interpretability.
@@ 3.1 Problem Definition
3.2 Overall Workflow
The CAF-I inference workflow (Figure 2) unfolds in key stages: Context (CA), Semantic (SA), and Rhetoric (RA) agents first independently provide initial assessments and reasoning. These are shared for collaborative reanalysis where agents refine judgments using peer insights. A Decision Agent (DA) then aggregates these refined judgments for an initial irony classification. Finally, a Refinement Evaluator Agent (REAgent) reviews this, potentially triggering a single conditional refinement loop with targeted feedback to improve the final classification within one iteration. The detailed design is presented in 3.3.
@@ 3.3.3 Rhetoric Agent
$D_{r}=\text{RhetoricalDeviceDetector}_{\text{LLM}}(x_{i})=\{d_{r1},...,d_{rm}\}$ (7)
@@ 4.1 Experiment Setup
To evaluate the efficacy and generalizability of CAF-I, we utilized four established sarcasm detection benchmarks: IAC-V1, IAC-V2, MUSTARD, and SemEval-2018 Task 3. Model performance was assessed using Accuracy and the Macro-F1 score, with Macro-F1 serving as the primary evaluation metric. The GPT-40 model was employed as the LLM backbone for all agents within the CAF-I framework, and a temperature setting of 0 was used for all API interactions to ensure reproducibility. We compared CAF-I against a comprehensive suite of baselines, including LLM-based approaches, fine-tuned Pre-trained Language Models, and deep learning method. For complete details regarding the experimental setup, including specific dataset characteristics, a detailed description of evaluation metrics, implementation specifics, and a comprehensive list of comparison baselines, please refer to Appendix A.
@@ 4.2 Overall Performance Comparison
MIARN (Tay et al., 2018) 63.21 63.18 72.75 72.75 64.60 63.90 68.50 67.80 67.26 66.91
SAWS (Pan et al., 2020) 66.13 65.60 76.20 76.20 69.71 70.95 69.90 68.90 70.48 70.41
DC-Net (Liu et al., 2021) 66.50 66.40 78.00 77.90 71.28 71.43 70.80 69.60 71.64 71.33
BERT (Devlin et al., 2019) 65.30 65.20 76.40 76.20 64.30 64.30 69.90 68.40 68.97 68.52
RoBERTa (Liu et al., 2019) 70.10 69.90 76.60 76.70 66.10 66.00 70.20 69.10 70.75 70.42
GPT-4o (Yao et al., 2024) 70.63 70.05 73.03 71.99 67.24 65.79 64.03 63.17 68.73 67.75
GPT-4o+CoC (Yao et al., 2024) 72.19 71.52 73.36 72.31 69.42 68.48 70.79 70.60 71.44 70.73
GPT-4o+Goc(Yao et al., 2024) 65.00 62.91 64.97 61.30 70.69 69.91 74.03 74.02 68.67 67.04
GPT-4o+Boc(Yao et al., 2024) 68.75 67.36 71.35 69.39 69.42 68.45 62.12 61.85 67.91 66.76
@@ 4.3 Ablation Study
To assess the contribution of each core component within CAF-I, we conducted ablation studies, with detailed results presented in Figure 3. These studies involved systematically removing specialized analysis agents CA, SA, or RA from all analysis stages, or deactivating the REAgent’s conditional refinement mechanism, thereby making the initial aggregated decision final. Experiments were performed on the IAC-V1, MuSTARD, and SemEval-2018 datasets.
@@ 4.4 Robust Experiment
To assess CAF-I’s architectural robustness, we evaluated its performance using alternative LLM backbones, specifically Qwen 2-7B and Llama 3-8B, on the four benchmark datasets while maintaining the core framework and prompts. These results were compared against each LLM’s baseline IO prompt performance, sourced from SarcasmCue (Yao et al., 2024), to isolate architectural contributions. The evaluation confirmed CAF-I’s architectural value. Across different LLM backbones, CAF-I consistently demonstrated substantial average Macro-F1 improvements over their respective IO prompting baselines. For instance, CAF-I with Qwen 2-7B achieved a notable 81.25% Macro-F1 on SemEval-2018, surpassing even its primary GPT-4o backbone implementation 79.99% Ma-F1 on this dataset. This general trend of outperformance strongly confirms CAF-I’s robustness and broad applicability. Detailed comparative results are presented in Appendix C.
@@ 4.5.1 Case Study
To illustrate the interpretability and error correction capabilities of CAF-I, a case study was conducted. In this instance (detailes are provided in Figure 4), CAF-I initially an incorrect ”Sarcasm” judgment arose because the Semantic and Rhetoric Agents over-interpreted certain phrases, despite the Context Agent finding no significant inconsistency. However, the REAgent identified this discrepancy and triggered a conditional refinement. In the subsequent feedback-guided re-evaluation, the Rhetoric Agents revised their assessments, aligning with the Context Agent towards ”Not Sarcastic.” This revised consensus allowed the Decision Agent to correctly classify the input, demonstrating REAgent’s effectiveness in error correction and enhancing judgment robustness.
@@ 4.5.2 Quantitative Evaluation of Explanations
Furthermore, to quantitatively assess the utility of the intermediate reasoning generated by CAF-I’s specialized agents, an auxiliary experiment is conducted. This evaluation, detailed in Appendix D, confirmed that the agent-generated explanations provide valuable discriminative information, thereby enhancing the framework’s interpretability.
@@ 5 CONCLUSION
Appendix A Detail Experiment Set Up
A.1 Datasets
  • IAC-V1 (Lukin & Walker, 2013): This dataset comprises comments from online political debates, offering examples of sarcasm embedded within argumentative discourse.
  • IAC-V2 (Oraby et al., 2017): An extension of IAC-V1, providing a larger, more diverse collection of sarcastic/non-sarcastic statements from similar online discussion forums.
  • MuSTARD (Castro et al., 2019): Sourced from popular television shows, MuSTARD provides conversational context for each utterance.
  • SemEval-2018 Task 3 (Van Hee et al., 2018): Consists of English tweets annotated for irony, representing challenges of social media text like brevity, informal language, and implicit contextual cues for ironic intent.
Dataset Year Size Avg. Length Domain Context
IAC-V1 2013 320 68 Debate No
IAC-V2 2016 1042 43 Debate No
MuSTARD 2019 784 14 Dialogue Yes
SemEval-2018 2018 183 14 Twitter No
A.2 Evaluation Metrics
We evaluate model performance using Accuracy for overall correctness and the Macro-F1 score as the primary metric, following standard practices in the field (Yao et al., 2024). Macro-F1 averages the F1-scores for ironic and non-ironic classes independently, providing a balanced measure suitable for potentially imbalanced datasets common in irony detection.
A.3 Experiment Details
A.4 Comparison Baselines
To contextualize CAF-I’s performance, we compare it against a comprehensive suite of baselines across three categories: LLM-based: Approaches from the SarcasmCue framework (Yao et al., 2024), including GPT-4o Zero-Shot and three advanced prompting strategies: Chain of Contradiction, Graph of Cues, and Bagging of Cues. Performance figures are cited from their reported results. Fine-tuned PLMs: Standard Pre-trained Language Models such as BERT-base-uncased (Devlin et al., 2019) and RoBERTa-base (Liu et al., 2019), fine-tuned on target datasets. Deep Learning Methods: Influential deep learning methods with task-specific architectures, namely MIARN (Tay et al., 2018), SAWS (Pan et al., 2020), and DC-Net (Liu et al., 2021). Results are sourced from literature with aligned evaluation settings (Qiu et al., 2024; Zhang et al., 2024; Tay et al., 2018; Xue et al., 2024; Pan et al., 2020).
Appendix B Case Study Illustration
This appendix section presents a detailed visual walkthrough of the case study discussed in Section 4.5.1. Figure 4 below illustrates the step-by-step multi-agent reasoning, collaborative reanalysis, and conditional refinement process undertaken by the CAF-I framework on an example input. This visualization aims to further clarify CAF-I’s interpretability and its mechanism for error correction, as highlighted in the main text, by showing how the agents interact and revise their judgments to reach a final, more accurate classification.
Appendix C Detailed Robustness Evaluation with Different LLM Backbones
This section provides the detailed performance metrics for the CAF-I framework when implemented with alternative Large Language Model backbones, namely Qwen 2-7B and Llama 3-8B. The results, as summarized in Table 4, are compared against each LLM’s baseline Input-Output prompt performance. This data supports the discussion in Section 4.4 regarding CAF-I’s architectural robustness and its ability to enhance performance across different underlying models.
Table 4: Robustness evaluation of the CAF-I framework using different LLM backbones Qwen 2-7B, Llama 3-8B compared to the IO prompting. Baseline IO results are aligned with SarcasmCue (Yao et al., 2024). All scores are in %.
LLM Backbone IAC-V1 IAC-V2 SemEval-2018 MuSTARD
Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
Qwen2-7b + IO 56.56 49.32 51.85 38.57 45.15 38.83 54.78 46.17
Qwen2-7b + CAF-I 71.85 71.19 67.08 66.39 81.63 81.25 72.60 73.80
Llama3-8b + IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99
Llama3-8b + CAF-I 60.62 60.40 70.36 73.55 75.76 75.47 59.56 57.11
Appendix D Quantitative Evaluation of Agent Explanations
Appendix E Future Work
Future research could focus on extending CAF-I’s application beyond irony detection to include other forms of emotional detection and broader nuanced linguistic phenomena, as well as adapting it for cross-lingual applications. Further work might also involve developing the framework for multimodal contexts, enhancing agent collaboration dynamics, and systematically evaluating its performance and efficiency with a wider array of Large Language Model backbones.

### v2

@@ Abstract
Keywords:
Irony Detection Multi-Agent Systems Large Language Model
@@ 1 Introduction
The rapid development of social media has fostered increasingly diverse and semantically complex forms of textual expression [9] highlights the challenge of irony detection. Irony, a rhetorical device contrasting stated and intended meaning [3], is crucial for NLP systems as its presence alters textual interpretation, impacting applications like sentiment analysis and content moderation [22]. However, detection is complicated by context dependency, obscure authorial intent, and varied rhetorical strategies [25].
Initial efforts used traditional machine learning with handcrafted features [22], often failing to capture irony’s subtleties. Deep learning methods improved semantic representation but struggled with the implicit and contextual nature of ironic utterances [8, 1]. Consequently, the emergence of LLM has presented novel avenues, exhibiting substantial potential in irony detection, typically via fine-tuning or advanced prompt engineering techniques [30].
@@
Motivated by Chain-of-Thought (CoT) prompting [26] and multi-agent cooperation principles [12], this paper introduces CAF-I, a novel LLM-driven multi-agent framework to address existing limitations in irony detection by simulating human-like multidimensional analytical processes. CAF-I initially performs multi-perspective parallel analysis from three core angles: contextual consistency, deep semantic logic, and rhetorical function. Subsequently, an interactive collaborative optimization mechanism facilitates information sharing and debate among these perspectives. Finally, an integrated decision mechanism allows a central decision-making agent to consolidate these viewpoints and apply conditional optimization triggered by evaluative feedback, forming the paper’s primary contributions:
@@ 2.1 Irony Detection
Irony detection in NLP remains challenging due to its inherent complexity, context dependence, and contradictory nature. Early machine learning relied on hand-crafted features, including lexical cues [5] and sentiment lexicons [22], but inadequately captured irony’s subtleties. Deep learning techniques offered significant improvements: Word embeddings such as Word2Vec [16] and GloVe [19] enabled richer semantic representations. Subsequently, Convolutional Neural Networks [20], recurrent neural networks, and particularly LSTM architectures [31] learned hierarchical and sequential features, while Graph Neural Networks [11] captured structural dependencies to improve performance. Nonetheless, these deep learning approaches still faced difficulties with irony’s implicit meanings and complex rhetoric, motivating the exploration of more advanced models.
Recent advances in large language models have introduced prompt-based and zero-shot learning methods, allowing models like GPT-4o to interpret irony with minimal reliance on labeled data [30]. By leveraging task-specific prompts, these models have demonstrated enhanced context-aware irony classification. However, existing LLM-based approaches still operate as single-model predictors, lacking the capacity for multi-perspective reasoning.
@@ 2.2 CoT Prompt Engineering
To address the limitations of direct predictions in complex reasoning tasks, CoT prompting emerged as a structured technique that enhances language models by guiding them through intermediate reasoning steps [26]. Despite its effectiveness, CoT initially depended on manually created prompts, which constrained its adaptability. Subsequent innovations, such as Auto-CoT, addressed this by automating the creation of reasoning chains [33]. Further extensions, such as Tree-of-Thought (ToT), expanded reasoning by enabling the exploration of multiple pathways [30], while Graph-of-Thought (GoT) introduced a structured approach where reasoning steps are represented as interconnected nodes in a graph [2]. These advancements directly influenced our multi-agent framework, where specialized agents leverage structured reasoning to collaboratively enhance irony detection.
@@ 2.3 Multi-agent Cooperation
Multi-agent frameworks effectively leverage collaborative interactions among specialized agents for complex tasks, with studies exploring forms like deliberation, structured debates, and dialogues. Debating frameworks, for instance, improve factual accuracy and solution diversity in complex reasoning [7]. Others, such as CAMEL [10], use role-play to simulate nuanced human behaviors; MathChat [28] employs structured dialogues for intricate tasks; and AutoGen [27] shows multi-agent versatility through customizable structures. Inspired by these methodologies, our framework uses specialized LLM-based agents for contextual, semantic, and rhetorical analysis via CoT reasoning. A final decision agent integrates their outputs, enhancing irony detection’s robustness, accuracy, and interpretability.
@@ 3.1 Problem Definition
3.2 Overall Workflow
The CAF-I inference workflow, depicted in Figure 2, unfolds in key stages. Initially, three specialized analysis agents—Context (CA), Semantic (SA), and Rhetoric (RA)—independently provide first-round assessments and reasoning. These outputs are shared for a collaborative reanalysis, where agents refine their judgments considering peer insights. A Decision Agent (DA) then aggregates these refined multiperspective judgments into an initial irony classification. Finally, a Refinement Evaluator Agent (REAgent) reviews this output and, if needed, triggers a single conditional refinement loop with targeted feedback to the analysis agents, aiming to improve the final classification within at most one iteration. The detailed design of each agent mentioned above is presented in 3.3.
@@ 3.3.3 Rhetoric Agent
$D_{r}=\text{RhetoricalDeviceDetector}_{\text{LLM}}(x_{i})=\{d_{r1},d_{r2},...,d_{rm}\}$ (7)
@@ 4.1 Experiment Setup
4.1.1 Datasets
  • IAC-V1 [15]: This dataset comprises comments from online political debates, offering examples of sarcasm embedded within argumentative discourse.
  • IAC-V2 [17]: An extension of IAC-V1, providing a larger, more diverse collection of sarcastic/non-sarcastic statements from similar online discussion forums.
  • MuSTARD [4]: Sourced from popular television shows, MuSTARD provides conversational context for each utterance.
  • SemEval-2018 Task 3 [24]: Consists of English tweets annotated for irony, representing challenges of social media text like brevity, informal language, and implicit contextual cues for ironic intent.
Dataset Year Size Avg. Length Domain Context
IAC-V1 2013 320 68 Debate No
IAC-V2 2016 1042 43 Debate No
MuSTARD 2019 784 14 Dialogue Yes
SemEval-2018 2018 183 14 Twitter No
4.1.2 Evaluation Metrics
We evaluate model performance using Accuracy for overall correctness and the Macro-F1 score as the primary metric, following standard practices in the field [30]. Macro-F1 averages the F1-scores for ironic and non-ironic classes independently, providing a balanced measure suitable for potentially imbalanced datasets common in irony detection.
4.1.3 Experiment Details
4.1.4 Comparison Baselines
To contextualize CAF-I’s performance, we compare it against a comprehensive suite of baselines across three categories: LLM-based: Approaches from the SarcasmCue framework [30], including GPT-4o Zero-Shot and three advanced prompting strategies: Chain of Contradiction, Graph of Cues, and Bagging of Cues. Performance figures are cited from their reported results. Fine-tuned PLMs: Standard Pre-trained Language Models such as BERT-base-uncased [6] and RoBERTa-base [13], fine-tuned on target datasets. Deep Learning Methods: Influential deep learning methods with task-specific architectures, namely MIARN [23], SAWS [18], and DC-Net [14]. Results are sourced from literature with aligned evaluation settings [21, 32, 23, 29, 18].
@@ 4.2 Overall Performance Comparison
MIARN [23] 63.21 63.18 72.75 72.75 64.60 63.90 68.50 67.80 67.26 66.91
SAWS [18] 66.13 65.60 76.20 76.20 69.71 70.95 69.90 68.90 70.48 70.41
DC-Net [14] 66.50 66.40 78.00 77.90 71.28 71.43 70.80 69.60 71.64 71.33
BERT [6] 65.30 65.20 76.40 76.20 64.30 64.30 69.90 68.40 68.97 68.52
RoBERTa [13] 70.10 69.90 76.60 76.70 66.10 66.00 70.20 69.10 70.75 70.42
GPT-4o [30] 70.63 70.05 73.03 71.99 67.24 65.79 64.03 63.17 68.73 67.75
GPT-4o+CoC [30] 72.19 71.52 73.36 72.31 69.42 68.48 70.79 70.60 71.44 70.73
GPT-4o+Goc[30] 65.00 62.91 64.97 61.30 70.69 69.91 74.03 74.02 68.67 67.04
GPT-4o+Boc[30] 68.75 67.36 71.35 69.39 69.42 68.45 62.12 61.85 67.91 66.76
@@ 4.3 Ablation Study
To quantitatively assess the contribution of each core component within CAF-I, we conducted ablation studies, with detailed results presented in Figure 3. These studies involved systematically removing specialized analysis agents CA, SA, or RA from all analysis stages, or deactivating the REAgent’s conditional refinement mechanism, thereby making the initial aggregated decision final. Experiments were performed on the IAC-V1, MuSTARD, and SemEval-2018 datasets.
@@
Table 3: Robustness evaluation of the CAF-I framework using different LLM backbones Qwen 2-7B, Llama 3-8B compared to the IO prompting. Baseline IO results are aligned with SarcasmCue [30]. All scores are in %.
LLM Backbone IAC-V1 IAC-V2 SemEval-2018 MuSTARD
Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
Qwen2-7b + IO 56.56 49.32 51.85 38.57 45.15 38.83 54.78 46.17
Qwen2-7b + CAF-I 71.85 71.19 67.08 66.39 81.63 81.25 72.60 73.80
Llama3-8b + IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99
Llama3-8b + CAF-I 60.62 60.40 70.36 73.55 75.76 75.47 59.56 57.11
@@ 4.4 Robust Experiment
To assess CAF-I’s architectural robustness beyond its GPT-4o backbone, we tested it with Qwen 2-7B and Llama 3-8B on four benchmark datasets, maintaining the core framework and prompts. Performance was compared against each LLM’s baseline IO prompt results, sourced from SarcasmCue [30], to isolate architectural contributions. Detailed results are in Table 3.
The evaluation validates CAF-I’s architectural value. Using alternative LLM backbones, CAF-I consistently yielded substantial average Macro-F1 improvements over their respective IO prompting baselines. Notably, on SemEval-2018, CAF-I with Qwen 2-7B achieved 81.25% Ma-F1, even outperforming its primary GPT-4o backbone implementation which scored 79.99% Ma-F1. This observation, alongside the general trend of outperforming IO baselines with different LLM, strongly confirms CAF-I’s robustness and broad applicability.
@@ 4.5.1 Case Study
Initially, an incorrect "Sarcasm" judgment arose because the Semantic and Rhetoric Agents over-interpreted certain phrases, despite the Context Agent finding no significant inconsistency. However, the REAgent identified this discrepancy and triggered a conditional refinement. In the subsequent feedback-guided re-evaluation, the Rhetoric Agents revised their assessments, aligning with the Context Agent towards "Not Sarcastic." This revised consensus allowed the Decision Agent to correctly classify the input, demonstrating REAgent’s effectiveness in error correction and enhancing judgment robustness.